OpenStudy (anonymous):

Suppose that the equation, y = 8(1.07)x, represents the winning prize money, in thousands, for first place in the a major auto race. X represents the number of years after 1920, the first year. What is the value for 1920 (x = 0)?

OpenStudy (anonymous):

Y = 8(1.07)^x X = 0 (1.07)^0 = 1 8(1) = 8 1920 = 8 thousand dollars.